# From Commands to Prompts: LLM-based Semantic File System for AIOS

### v1

@@ Abstract
Large language models (LLMs) have demonstrated significant potential in the development of intelligent applications and systems such as LLM-based agents and agent operating systems (AIOS). However, when these applications and systems interact with the underlying file system, the file system still remains the traditional paradigm: reliant on manual navigation through precise commands. This paradigm poses a bottleneck to the usability of these systems as users are required to navigate complex folder hierarchies and remember cryptic file names. To address this limitation, we propose an LLM-based semantic file system (LSFS) for prompt-driven file management. Unlike conventional approaches, LSFS incorporates LLMs to enable users or agents to interact with files through natural language prompts, facilitating semantic file management. At the macro-level, we develop a comprehensive API set to achieve semantic file management functionalities, such as semantic file retrieval, file update monitoring and summarization, and semantic file rollback). At the micro-level, we store files by constructing semantic indexes for them, design and implement syscalls of different semantic operations (e.g., CRUD, group by, join) powered by vector database. Our experiments show that LSFS offers significant improvements over traditional file systems in terms of user convenience, the diversity of supported functions, and the accuracy and efficiency of file operations. Additionally, with the integration of LLM, our system enables more intelligent file management tasks, such as content summarization and version comparison, further enhancing its capabilities. The code is open-source at https://github.com/agiresearch/AIOS-LSFS.
@@ 1 Introduction
In recent years, researchers have integrated AI with operating systems to create innovative, efficiency-boosting applications mei2024aios ; blair1987knowledge . For example, machine learning algorithms can dynamically adjust system resource allocation schneider2020machine ; gong2024dynamic . Besides, by monitoring application performance and user behavior, AI models can predict future resource needs and pre-allocate CPU, memory, and other resources, enhancing overall system performance and response speed KAUR2023101804 . With the recent developments of generative AI, researchers have envisioned Large Language Models (LLMs) based operating systems ge2023llm .
At present, LLMs have demonstrated great planning and reasoning ability yao2022react ; wei2022chain ; hao2023reasoning ; zhao2024large ; jin2024impact , which facilitates the development of LLM-based agents, including single-agent applications yang2024swe ; zhang2023you ; gur2023real ; deng2024mind2web and collaborative multi-agent applications wu2024autogen ; ge2024openagi ; shen2024hugginggpt ; hong2023metagpt . Nowadays, researchers have been trying to embed LLMs from the application-level to the system-level to serve users and AI agent applications, enabling LLM-based artificial intelligent operating system (AIOS) mei2024aios . Through such integration, the powerful understanding and generation ability of LLMs not only can help to complete tasks on the application-level, but also can help to intelligently manage various resources such as memory and storage on the system-level.
File system is one of the most fundamental and essential components of an operating system, yet it comes with significant limitations in terms of AIOS. Currently, file management primarily relies on index structures that organize files based on attributes like file name, size, and timestamps. However, this approach does not leverage the semantic meaning of the files themselves. For instance, if two files have similar content–such as different versions of the same document–traditional file systems lack the capability to organize or retrieve these files based on their content similarity. User interactions with traditional file systems require complex operating system commands or manual navigation through the user interface, forcing users to precisely recall file names or locations. For systems with numerous files, this retrieval process can be inefficient and time-consuming, reducing overall system usability. Nowadays, based on the strong language understanding capability of LLMs, we can make better use of the file content and semantic information for file management by introducing LLMs into the system. However, existing works on using LLMs to facilitate file management are mostly conducted on the application level, which targets at designing specific agent for file retrieval and manipulation liu2024agentlite ; talebirad2023multi . The community still lacks a more general LSFS to serve as a common foundation that can be used by various agents on the application-level. To solve the above problems, we propose LSFS, an LLM-based semantic file system, to enhance the interaction between LLM and traditional file system for better semantic file management.
@@
To address problem , we integrate LLM into the API for complex functions and introduce a template system prompt. This allows us to utilize LLM to extract keywords from the user’s natural language input and map them directly as input parameters for the API or syscall, streamlining the interaction between users and the system. The comparison of commands executed by users in traditional file systems and LSFS is shown in  Figure 1. In  Figure 1(a), When a user wishes to modify a file’s content, they must input a specific command in the terminal, requiring them to remember the correct operator and the exact paths for both the target and source files. This approach places a significant burden on the user’s memory. However, LSFS can effectively solve this problem. Users only need to manage files by typing a natural language prompt as a command. For example, as shown in Figure Figure 1(b), the user needs to input a simple natural language description, our LSFS is able to understand prompts and perform the corresponding operation. This approach greatly simplifies the complexity of the user’s operation and improves the efficiency of the user’s operation. Furthermore, to reduce the LLM’s hallucination problem of generating inaccurate instructions, especially those irreversible operations, we design systematic safety insurance mechanisms in LSFS, such as safety checks for irreversible operations and user verification before instruction execution.
@@
  • Through our experiments, we validate the completeness of LSFS’s functions, while also demonstrating its innovation and superiority in handling specific tasks compared to traditional file systems. Furthermore, the experiments suggest that the LSFS enhances the execution of file management operations, making it more intelligent and efficient.
@@ 2.1 Semantic File System
Currently, file storage and retrieval primarily rely on an index structure maintained by the system, where file metadata points to the file’s location on the disk dai2022state . While optimizing the index structure can enhance retrieval efficiency, the current storage model is still largely dependent on the keywords extracted from the file’s content. Gifford et al. gifford1991semantic were the first to propose a semantic file system, which introduced a layer that generates directories by extracting attributes from files, enabling users to query file attributes through navigation. eck2011semantic proposed a semantic file system to manage the data. Many subsequent works have integrated semantics into metadata hua2011semantic ; mahalingam2003towards ; hua2009smartstore ; hardy1993essence ; mohan2006semantic . Mahalingam et al. mahalingam2003towards proposed a semantic deep file management system called Sedar, which embeds semantic features into file metadata, allowing retrieval to be matched with semantic vectors. However, Sedar faces challenges with the similarity of extracted semantic vectors. Hua et al. hua2011semantic further integrated semantic information into metadata, restricting complex queries to a related metadata group. The same authors hua2013sane also utilized semantics to reduce the relevance of queries using the semantic similarity between files based on the semantic naming system. Leung et al. leung2009copernicus used semantic information combined with the file system design of graphs to provide scalable search and navigation. On the other hand, Bloehdorn et al. bloehdorn2006tagfs proposed to manage files through semantic tags. Schandl et al. schandl2009sile developed an approach for managing desktop data using a semantic vocabulary. In contrast, our semantic file system is based on the strong language understanding ability of LLMs. Besides, it integrates comprehensive semantic information across all aspects of the file system–from storage and file operations to practical applications. This holistic approach enhances the system’s ability to understand and manage files, significantly improving functionality beyond what traditional systems and earlier semantic file systems offer.
@@ 2.2 Semantic Parser
Researchers have also devoted efforts to developing semantic parsers kamath2018survey ; mooney2007learning ; wong2006learning ; clarke2010driving ; yih2014semantic capable of transforming natural language into a machine-interpretable format. Mooney et al. mooney2007learning pioneered this area by introducing a suite of machine learning-based semantic parsers designed for specific tasks. Iyer et al. iyer2017learning subsequently focused on parsing database commands, while Berant et al. berant2013semantic proposed a question-answer pair learning approach to enhance parsing efficiency. In further work, the same authors explored a paraphrasing technique berant2014semantic to improve semantic parsing performance. Poon et al. poon2009unsupervised introduced a Markov logic-based approach, and Wang et al. wang2015building addressed the challenge of building parsers from scratch in new domains. Ge et al. ge2005statistical proposed a parse tree-based method for more accurate semantic analysis. Notably, Lin et al. lin2018nl2bash were the first to integrate a semantic parser into an operating system, leveraging a dataset of bash commands and expert-written natural language to establish a mapping between the two. However, this approach faced limitations in handling complex semantics and unseen natural language. In contrast, our LSFS is built upon more constructing more semantic index as the format of embedding vectors, improving its generation ability to understand and process diverse natural language inputs.
2.3 OS-related LLMs and Agents
To enhance user experience and improve task completion efficiency, many LLM-based AI agents have been introduced to fulfill operating system related tasks such as software development yang2024swe ; qian2023communicative and manipulating resources on traditional OS wu2024copilot ; bonatti2024windows . Wu et al. wu2024copilot developed LLM-based agents for co-piloting users’ interaction with computers, such as drawing charts and creating web pages. MetaGPT hong2023metagpt employs a sophisticated large language model in a multi-agent conversational setting to automate software development, assigning specific roles to various GPTs for seamless collaboration. CHATDEV qian2023communicative presents a framework for software development, utilizing agents to enhance cooperation among different roles within the development process. Beyond the application-level research on LLM-based agents, researchers also explored integrating LLMs into the system-level mei2024aios ; bonatti2024windows , which enables LLM-based agent operating system (AIOS) that provides fundamental services to the agent applications on top of it, such as agent scheduling, memory and storage management, tool management, security features, and access control.
@@ 3 Architecture of LSFS
This system retains traditional file attributes and integrates semantic information derived from the file’s content. Unlike traditional file systems, our system enhances metadata by including user-defined keywords that capture the essence of the file’s content, in addition to standard file attributes. Our semantic file management system employs embedding models to convert textual content into semantic vector representations. These vectors enable the construction of a semantic-based data index, facilitating the structured storage of data and enabling robust, semantically-driven file retrieval and management. This optimized architecture allows for more intelligent and semantic interactions with file data, enhancing the system’s functionality and making it particularly well-suited for modern file management tasks.
Figure 2(a) outlines the overall architecture of our LSFS. LSFS operates as an additional layer on top of traditional file systems, which constructs semantic index of the file in the format of embedding vectors, thereby enabling more advanced content-based file management operations. LSFS provides a range of system calls (syscalls) that support complex file operations and reuse. The system also includes a supervisor that works in tandem with the traditional file management system. This supervisor continuously monitors changes in the traditional file system and synchronizes those changes in LSFS, ensuring that the contents of LSFS remain consistent with the underlying file system.  Figure 2(b) presents an overview of the syscall structure in LSFS. These syscalls are primarily categorized into two types: atomic syscalls and composite syscalls. The former handles basic operations such as adding, deleting, modifying, and searching files, while the latter builds upon atomic syscalls to execute more complex functions. On the top of the syscalls, LSFS APIs are constructed by utilizing one or more syscalls and integrating LLMs. These APIs facilitate more advanced yet general-purpose file operations, demonstrating the superiority of LSFS in handling such tasks. Further details on these capabilities will be discussed in Section 4.3.
@@ 4 Implementation of LSFS
Function TFS LSFS (Ours) create new directory mkdir() create_or_get_file() open file open() create_or_get_file() read file read() create_or_get_file() get file state and metadata stat() create_or_get_file() delete directory rmdir() del_() delete file unlink() / remove() del_() write data write() insert() overwrite data write() overwrite() update the access time utime() update_access_time() automatic comparison compare_change() generate link symlink() / link() / readlink() generate_link() lock or unlock file flock() lock_file() / unlock_file() rollback snapshot + rollback rollback() file translation file_translate() file group group_text() / group_se() merge file cat join() keyword retrieve grep keyword_retrieve() semantic retrieve semantic_retrieve() Hybrid retrieval integrated_retrieve()
We introduce the implementation of LSFS from the bottom to the top in the LSFS architecture shown in the Figure 2. We start by introducing the basic syscalls implemented in LSFS and introduce the supervisor which interacts between LSFS syscalls and traditional file systems. Then we present the APIs that built upon the syscalls to achieve more complex functionalities. After that, we introduce the LSFS parser on top to show how natural language prompts have been decoded into executable LSFS APIs. At last, we use different concrete prompts to show how different modules in the LSFS are executed to achieve functionalities.
@@ 4.2 Supervisor
The supervisor is implemented to track the changes in the files in the disk and sync the changes to the LSFS. The supervisor periodically scans the files within its specified directory. When it detects any change or deletion of the file content, it automatically synchronizes this information with the LSFS by invoking the appropriate syscall. This ensures that the state of the file in the LSFS reflects the current state of the file in memory. In order to ensure the accuracy of other syscall operations on the file after Supervisor modifies the internal file status of LSFS, LSFS introduces the process lock mechanism to ensure that multiple processes can access the file correctly. When asked, the supervisor can generate status change reports. For example, when a file is modified, the supervisor invokes LLMs to generate a detailed modification log, compares the contents of the file before and after modification, and presents this log to the user.
@@ 4.3 API of LSFS
In this module, we introduce the APIs that are implemented on top of the syscalls mentioned in Section  2 to support higher-level semantic file management functions. Specifically, we provide the following APIs that cover the basic semantic file management requirements, i.e., semantic CRUD (create, read, update, and delete) of files.
Retrieve-Summary API.
Change-Summary API.
Rollback API.
This API is designed to achieve version restoration by utilizing the version recorder from the change-summary API, making version rollbacks more manageable. In this API, the overwrite() and create_or_get_file() syscalls are employed. Traditional file systems like ZFS, BtrFS, and NTFS offer rollback capabilities, but they primarily rely on snapshots, where the system captures the state of files at a specific point in time and restores them to that version. Our rollback API in the LSFS introduces two rollback methods for greater flexibility and ease of use. The first is time-based rollback, where the LLMs parses the rollback time from the user’s input and reverts the file to the corresponding version. The second is version-based rollback, allowing users to specify in the prompt how many versions backward they wish to revert. This dual approach makes it easier for users to rollback to the target version of files.
Link API.
This API is designed to enable the creation of shareable file links. In traditional file systems, links can only be generated for local access, limiting collaboration. However, with the LSFS’s link API, shareable links can be generated for broader accessibility. Specifically, the API leverages cloud database, e.g., Google Drive, to upload files and generate the shareable link. Additionally, validity period can be passed as an argument for this API, once the period expires, the link API automatically revokes the link and terminates access for secure and time-bound file sharing.
@@ 4.4 LSFS Parser
We design an LSFS parser to guide this conversion process and create a general prompt template for future API designs to extract parameters from natural language prompts. As illustrated in the accompanying  Figure 3, we leverage large language model to parse natural language commands. For each API, we design distinct system prompt words, which, when input alongside the user’s command, allow the language model to automatically detect and extract the key parameters, outputting them in a comma-separated format. By applying a simple regex to this output, we can directly convert the extracted keywords into syscall parameters, enabling seamless execution of the API command. This allows us to efficiently map natural language commands to the exacted parameters required by the system calls, benefiting interactions between natural language prompts and LSFS.

### v5

@@ Abstract
[…]
@@ 1 Introduction
In recent years, researchers have put great efforts in integrating AI to provide better services for serving applications. For example, machine learning algorithms have been studied to optimize system resource allocation and and improve system efficiency (blair1987knowledge; schneider2020machine; gong2024dynamic). The emergence of large language models (LLMs) has further catalyzed the integration of AI into serving applications. The great reasoning and planning ability of LLMs facilitates the development of LLM-based agents, including single-agent applications (yang2024swe; zhang2023you; gur2023real; deng2024mind2web) and collaborative multi-agent applications (wu2024autogen; ge2024openagi; shen2024hugginggpt; hong2023metagpt).
In these LLM-based agents, file management operations primarily rely on the traditional way that builds index structures that organize files based on attributes like file name, size, and timestamps. However, this approach still relies on the traditional file system semantics (e.g., POSIX), which is unable to leverage the high-level semantic meaning in the natural language context. For instance, if two files have similar content which cannot be distinguished by simple string matching, traditional file systems lack the ability to organize or retrieve these files based on content similarity. User interactions with traditional file systems require complex operating system commands or manual navigation through the user interface, forcing users to precisely recall file names or locations. For systems with numerous files, this retrieval process can be inefficient and time-consuming, reducing overall system usability. Nowadays, based on the strong language understanding capability of LLMs, we can make better use of the file content and semantic information for file management by introducing LLMs into the system. However, existing works on using LLMs to facilitate file management are mostly conducted on the application level, which targets at designing specific agent for file retrieval and manipulation (liu2024agentlite; talebirad2023multi). The community still lacks a more general LSFS to serve as a common foundation that can be used by various agents on the application-level. To solve the above problems, we propose LLM-based Semantic File System (LSFS) to more effectively integrate LLM and traditional file system for semantic file management.
@@
To address problem , we integrate LLM into the API for complex functions and introduce a template system prompt. This allows us to utilize LLM to extract keywords from the natural language of user input and map them directly as input parameters for the API or syscall, streamlining the interaction between users and the system. The comparison of commands executed by users in traditional file systems and LSFS is shown in  Figure 1. In  Figure 1(a), When a user wishes to modify a contetn of file, they must input a specific command in the terminal, requiring them to remember the correct operator and the exact paths for both the target and source files. This approach places a significant burden on the user’s memory. However, LSFS can effectively solve this problem. Users only need to manage files by typing a natural language prompt as a command. For example, as shown in Figure 1(b), the user just needs to input a simple natural language description, our LSFS is able to understand prompts and perform the corresponding operation. This approach greatly simplifies the complexity of the operation of user and improves the efficiency of the operation of user. Furthermore, to reduce the hallucination problem of generating inaccurate instructions of LLM, especially those irreversible operations, we design systematic safety insurance mechanisms in LSFS, such as safety checks for irreversible operations and user verification before instruction execution.
@@
[…]
@@ 2.1 Semantic File System
Currently, file storage and retrieval primarily rely on an index structure maintained by the system, where file metadata points to the location of file on the disk (dai2022state). While optimizing the index structure can enhance retrieval efficiency, the current storage model is still largely dependent on the keywords extracted from the content of file. Gifford et al. (gifford1991semantic) were the first to propose a semantic file system, which introduced a layer that generates directories by extracting attributes from files, enabling users to query file attributes through navigation. (eck2011semantic) proposed a semantic file system to manage the data. Many subsequent works have integrated semantics into metadata (hua2011semantic; mahalingam2003towards; hua2009smartstore; hardy1993essence; mohan2006semantic). (hua2013sane) utilized semantics to reduce the relevance of queries using the semantic similarity between files based on the semantic naming system. Leung et al. (leung2009copernicus) used semantic information combined with the file system design of graphs to provide scalable search and navigation. On the other hand, Bloehdorn et al. (bloehdorn2006tagfs) proposed to manage files through semantic tags. Schandl et al. (schandl2009sile) developed an approach for managing desktop data using a semantic vocabulary. In contrast, our semantic file system is based on the strong language understanding ability of LLMs. Besides, it integrates comprehensive semantic information across all aspects of the file system–from storage and file operations to practical applications. This holistic approach enhances the ability of system to understand and manage files, significantly improving functionality beyond what traditional systems and earlier semantic file systems offer.
@@ 2.2 Semantic Parser
Researchers have also devoted efforts to developing semantic parsers (kamath2018survey; mooney2007learning; wong2006learning; clarke2010driving; yih2014semantic) capable of transforming natural language into a machine-interpretable format. Iyer et al. (iyer2017learning) subsequently focused on parsing database commands, while Berant et al. (berant2013semantic) proposed a question-answer pair learning approach to enhance parsing efficiency. In further work, the same authors explored a paraphrasing technique (berant2014semantic) to improve semantic parsing performance. Poon et al. (poon2009unsupervised) introduced a Markov logic-based approach, and Wang et al. (wang2015building) addressed the challenge of building parsers from scratch in new domains. Ge et al. (ge2005statistical) proposed a parse tree-based method for more accurate semantic analysis. Notably, Lin et al. (lin2018nl2bash) were the first to integrate a semantic parser into an operating system, leveraging a dataset of bash commands and expert-written natural language to establish a mapping between the two. However, this approach faced limitations in handling complex semantics and unseen natural language. In contrast, our LSFS is built upon constructing semantic indexes for files in the format of embedding vectors, improving its generation ability to understand and process diverse natural language inputs.
2.3 OS-related LLM-based Agents
Many works (yang2024swe; qian2023communicative; wu2024copilot; bonatti2024windows; wang2024mobile; wang2024mobile2; yang2023appagent) introduce LLM-based agents to solve OS-related tasks. To help users solve more practical OS-related tasks with natural language interaction, different agents are proposed for both PCs (wu2024copilot; bonatti2024windows) and mobile devices (wang2024mobile; wang2024mobile2; yang2023appagent). Wu et al. (wu2024copilot) developed LLM-based agents for co-piloting users’ interaction with computers, such as drawing charts and creating web pages. MetaGPT (hong2023metagpt) employs a sophisticated large language model in a multi-agent conversational setting to automate software development, assigning specific roles to various GPTs for seamless collaboration. CHATDEV (qian2023communicative) presents a framework for software development, utilizing agents to enhance cooperation among different roles within the development process. Beyond the application-level research on LLM-based agents, researchers also explored integrating LLMs into the system-level (mei2024aios; bonatti2024windows), which target at low-level and general management services (e.g., scheduling and resource allocation) for agent applications running on the top.
@@ 3 Architecture of LSFS
This system retains traditional file attributes and integrates semantic information derived from the content of file. Unlike traditional file systems, our system enhances metadata by including user-defined keywords that capture the essence of the content of file, in addition to standard file attributes. Our semantic file management system converts textual content into semantic vector representations. These vectors enable the construction of a semantic-based data index, facilitating the structured storage of data and enabling robust, semantically-driven file retrieval and management. This optimized architecture allows for more intelligent and semantic interactions with file data, enhancing the system’s functionality and making it particularly well-suited for modern file management tasks.
Overview of the architecture.
Figure 2(a) outlines the overall architecture of our LSFS. LSFS operates as an additional layer on top of traditional file systems, working as a bridge between agents/users and traditional file systems. In the LSFS, we leverage an embedding model, i.e., all-MiniLM-L6-v2 (reimers-2019-sentence-bert), which is commonly used in vector database, to build index for files in the format of embedding vectors, thereby supporting more advanced file operations which requires semantic understanding of file content. LSFS provides a range of system calls (syscalls) that support complex file operations and reuse. The system also includes a supervisor that works in tandem with the traditional file management system. This supervisor continuously monitors changes in the traditional file system and synchronizes those changes in LSFS, ensuring that the contents of LSFS remain consistent with the underlying file system.  Figure 2(b) presents an overview of the syscall structure in LSFS. These syscalls are primarily categorized into two types: atomic syscalls and composite syscalls. The former handles basic operations such as adding, deleting, modifying, and searching files, while the latter builds upon atomic syscalls to execute more complex functions. On the top of the syscalls, LSFS APIs are constructed by utilizing one or more syscalls and integrating LLMs. These APIs facilitate more advanced yet general-purpose file operations, enhancing the capability of LSFS to solve file operation tasks require more semantic understandings. The implementations of these advanced APIs will be discussed in Section 4.3.
@@ 4 Implementation of LSFS
Function Implementation in TFS Implementation in LSFS create new directory mkdir() create() create file touch() create_or_get_file() open file open() create_or_get_file() read file read() create_or_get_file() get file state and metadata stat() create_or_get_file() delete directory rmdir() del_() delete file unlink() / remove() del_() write data write() add_() overwrite data write() overwrite() update the access time utime() update_access_time() automatic comparison compare_change() generate link symlink() / link() / readlink() generate_link() lock or unlock file flock() lock_file() / unlock_file() rollback snapshot + rollback rollback() file group group_keywords() / group_semantic() merge file cat file_join() keyword retrieve grep keyword_retrieve() semantic retrieve semantic_retrieve() Hybrid retrieval integrated_retrieve()
We introduce the implementation of LSFS from the bottom to the top in the LSFS architecture shown in the Figure 2. In the following parts, we start by introducing the basic syscalls implemented in LSFS and introduce the supervisor which interacts between LSFS syscalls and traditional file systems. Then we present the APIs that built upon the syscalls to achieve more complex functionalities. After that, we introduce the LSFS parser on top to show how natural language prompts have been decoded into executable LSFS APIs. At last, we use different concrete prompts to show how different modules in the LSFS are executed to achieve functionalities.
@@ 4.2 Supervisor
The supervisor is implemented to track the changes in the files in the disk and sync the changes to the LSFS. The supervisor periodically scans the files within its specified directory. When it detects any change or deletion of the file content, it automatically synchronizes this information with the LSFS by invoking the appropriate syscall. This ensures that the state of the file in the LSFS reflects the current state of the file in memory. LSFS also leverages the process lock mechanism to ensure that multiple processes can access the file correctly without sychronization problem. The supervisor also supports the change log, for example, when a file is modified, the supervisor invokes the LLM to generate a detailed modification log, compares the contents of the file before and after modification.
@@ 4.3 API of LSFS
In this module, we introduce the APIs that are implemented on top of the syscalls mentioned in Section 4.1 to support higher-level semantic file management functions. Specifically, we provide the following APIs that cover the basic semantic file management requirements, i.e., semantic CRUD (create, read, update, and delete) of files. The details of APIs are presented in Appendix B.
  • Retrieve-Summary API. This API implements the retrieval operations in LSFS, including keyword search and semantic search, and feeds back the retrieved content to the user through LLM.
  • Change-Summary API. This API implements the modification of the file content of the object file in LSFS, and compares the contents of the file before and after modification through LLM and gives a summary.
  • Rollback API. This API allows you to rollback a given file and provides several ways to do so, which includes rollback by date or rollback by version number.
  • Link API. This API generates a shareable link for a given file. Users can set an expiration date for the link, after which the link will be invalid.
We also design and implement the LSFS parser to parse natural language prompts into API calls that can be executed in the LSFS, which will be introduced in Section 4.4.
@@ 4.4 LSFS Parser
To parse natural language prompts into executable API calls, we implement a LSFS parser based on the LLMs and designs well-structured json-format schemas for each API inside the parser.
@@
Without specific mention, our parser is based on GPT-4o-mini by default, evaluations of using other LLMs will be reports in Section 5.1. By leveraging this parser, natural language prompts can be parsed into executable API calls (i.e., API function names and API function arguments), enabling seamless execution of the API command. This can help deal with natural language prompts with multiple and complex situations, benefiting interactions between natural language prompts and LSFS. As illustrated in the accompanying  Figure 3, when input alongside the command of user, the LSFS parser is able to extract the key parameters, including function names and arguments in a comma-separated format.
4.5 The Interaction between Modules
In Figure 4, we present the examples to demonstrate how components of LSFS interact with each other to achieve different functionalities. The upper section of Figure 4 depicts the workflow of the retrieve-summary API, while the lower section outlines the workflows of the change-summary API and rollback API. In the upper part of Figure 4, the LSFS parser decodes prompts into API calls with API name and API arguments. Then LSFS executes the API to check vector database to get results. We used llamaindex to index the database and subsequently retrieved the contents of the vector database by llamaindex. This API also provides user-interaction interface for the users to verify results. After the verification, the content will be summarized by leveraging LLM. In the lower part of Figure 4, when a modification request is submitted, the LSFS parser decodes the file information (name and location) that is to be modified. The LSFS then modifies the semantic changes in both the vector database and the files stored in the disk. Meanwhile, the supervisor of the LSFS is kept running to ensure consistency between the semantic index of files in the LSFS and the files stored in the disk. Upon updated, the summarization API compares the file contents before and after the change to generate a detailed change log. Additionally, the API stores the pre-modification content in the version recorder. If a rollback is requested, the API retrieves the specific version from the version recorder and synchronizes it in both the LSFS and files in the disk.